OpenStudy (anonymous):

what is the first and second derivative of xe^(2x)?

OpenStudy (anonymous):

Remember the product rule? (fg)'=f'g+fg' You have to use it. Try with f=x and g=e^(2x) Do you know how to go on?

OpenStudy (anonymous):

not really, so it's not 2xe^2x?

OpenStudy (anonymous):

No, you have to use the product rule. Find this first: f=x ---> f'=? g=e^(2x) ---> g'=? What did you get?

OpenStudy (anonymous):

x^2 and 2e^2x

OpenStudy (anonymous):

x^(2)e^(2x) + 2xe^(2x)?

OpenStudy (anonymous):

f=x -------> f'=1 g=e^(2x)---->g'=2e^(2x) Plug that on this formula: f'g+fg'

OpenStudy (anonymous):

e^2x + 2xe^2x

OpenStudy (anonymous):

so the next derivative would be 2e^2x + 4xe^2x?

OpenStudy (lgbasallote):

\[\large xe^{2x}\] \[\Large y' = 2xe^{2x} + e^{2x}\] \[\LARGE y'' = 2(xe^{2x} + e^{2x}) + e^{2x} \implies 2xe^{2x} + 2e^{2x} +e^{2x}\] if my solution is right...that means your answer is wrong..

OpenStudy (anonymous):

i don't understand how you can pull the 2 out like that

OpenStudy (lgbasallote):

constant multiple rule...

OpenStudy (lgbasallote):

are you familiar with that rule?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

\[\huge \frac{d}{dx} (ax) \implies a [\frac{d}{dx} (x)]\] where a is a constant

OpenStudy (lgbasallote):

for example derivative of 2x^2 is 2( derivative of x^2) derivative of x^2 = 2x so 2(derivative of x^2) = 2(2x) = 4x

OpenStudy (lgbasallote):

i think you need to review the rules of derivatives http://www.math.brown.edu/UTRA/derivrules.html

OpenStudy (anonymous):

how the helll am i supposed to plug that into a Wronskian?!

OpenStudy (lgbasallote):

hmm wait...i missed something.. \[\Large 2(2xe^{2x} + e^{2x}) + e^{2x} \implies 4xe^{2x} + 2e^{2x} + e^{2x}\]

OpenStudy (lgbasallote):

now combine.. \[\large \implies 4xe^{2x} + 3e^{2x}\] now do your wronskian thingy

OpenStudy (anonymous):

this is bull.... when am i really going to use the wronksian and cramer when I am an actuarry?!

OpenStudy (anonymous):

this one problem is taking forever!